OpenStudy (anonymous):

Use mathematical induction to prove the statement is true for all positive integers n, or show why it is false. medal!!!!! @ganeshie8

OpenStudy (anonymous):

4 ⋅ 6 + 5 ⋅ 7 + 6 ⋅ 8 + ... + 4n( 4n + 2) =4(4n+1)(8n+7)/6

OpenStudy (anonymous):

@SithsAndGiggles do you know this?

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

I've done my share of induction proofs, yes. Give me a moment

OpenStudy (anonymous):

please show all steps so i understand it

OpenStudy (anonymous):

okay thanks(: @SithsAndGiggles

OpenStudy (anonymous):

What happens when \(n=1\)?

OpenStudy (anonymous):

i dont know:(

OpenStudy (anonymous):

What I'm asking is, "is this true?" \[(4\cdot1)(4\cdot1+2)=\frac{4(4\cdot1)((8\cdot1)+7)}{6}\]

OpenStudy (anonymous):

no it is false

OpenStudy (anonymous):

Okay, so if the formula above doesn't hold for \(n=1\), then it's certainly true that it does not hold for ALL natural numbers. What does that tell you? Is the formula correct?

OpenStudy (anonymous):

since it is false, it is not correct

OpenStudy (anonymous):

Right. The counter-example is \(n=1\), so the statement/formula is false.

OpenStudy (anonymous):

so wait since it is false all i have to do is write that it is false? it shows to prove that the statement is false... how to i do that?

OpenStudy (anonymous):

nevermind, i understand(: thanks!

OpenStudy (anonymous):

yw